OpenStudy (anonymous):

solve 2 sqrt x+7 - 10 = -8

OpenStudy (anonymous):

a) x = -6 b) x = -4 c) x = 32 d) x = 74

OpenStudy (anonymous):

Add 10 to both sides, then divide both sides by 2. Then square both sides. Then subtract 7 from both sides. What do you get?

OpenStudy (anonymous):

2 sqrt x+7 = 2

OpenStudy (anonymous):

The only way this equation makes sense is if it actually means:\[2 \sqrt{x + 7} - 10 = -8\]instead of what is written in the problem statement. So, if you follow those four steps from my first post, you will get your answer.

OpenStudy (anonymous):

You did the first step, now do the second step.

OpenStudy (anonymous):

\[\frac{ 2\sqrt{x + 7} }{ 2 } = \frac{ 2 }{ 2 } \times \sqrt{x + 7} = 1 \times \sqrt{x + 7}\]